OpenStudy (anonymous):

help please! Evaluate the series: Enter only a numerical answer. I'm not sure if you would add up 1^2...20^2 on this one?

OpenStudy (anonymous):

\[\sum_{k=1}^{20} k\]

OpenStudy (anonymous):

\[\sum_{k=1}^{20}k ^{2}=1^{2}+2^{2}+3^{2}+...+20^{2}\] \[\sum n ^{2}=\frac{ n \left( n+1 \right)\left( 2n+1 \right) }{6 }\] put n=20 and get the solution.

OpenStudy (anonymous):

2870?

OpenStudy (anonymous):

yes you are correct.